OpenStudy (anonymous):

Abhasra made a trip to the town hall and back.The trip there took six hours and the trip back took five hours. She averaged 8km/h faster on the return trip than on the outbound trip. What was Abhasra's average speed on the outbound trip? I go the answer v=6. Is this right

OpenStudy (anonymous):

just want to know if its right

OpenStudy (kropot72):

Let v = speed on outbound trip Then speed on return trip = v + 8 Distance = speed * time v * 6 = (v + 8) * 5 6v = 5v + 40 ...........(1) Now you need to solve equation (1) to find the value of v.

OpenStudy (anonymous):

so its wrong then

OpenStudy (anonymous):

Where did you get the one though ?

OpenStudy (kropot72):

Your answer v = 6 is not correct. "...............(1)" is just a way of identifying the equation to be solved, which is 6v = 5v + 40

OpenStudy (anonymous):

So its v=40?

OpenStudy (kropot72):

Yes, v = 40 mph.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

stay I may need another help

OpenStudy (kropot72):

Sorry, i must log out now. Please post your questions for others to help :)

OpenStudy (anonymous):

Darryl traveled to the lake and back. It took one hour longer to go there than it did to come back. The average speed on the trip there was 36 mph. The average speed on the way back was 45 mph. How many hours did the trip there take?